# Did I Faithfully Say What I Thought? Bridging the Gap Between Neural Activity and Self-Explanations in Large Language Models

### v1

@@ 1 Introduction
Autoregressive Transformer-based Vaswani et al. (2017) Large Language Models (LLM) have demonstrated striking performance in a wide range of Natural Language Processing (NLP) tasks Chang et al. (2024). This achievement was largely attributable to the reasoning LLM abilities, enabling them to break down complex multi-step problems with Chain-of-Thought (CoT) Wei et al. (2024) and generate self free text Natural Language Explanation (self-NLE) to support their answers Wiegreffe and Marasovic ; Huang et al. (2023). Generating CoT consists in prompting the LLM to generate an explanation in an explain-then-predict fashion, intending to improve the model’s performance. Conversely, generating self-NLE follows a predict-then-explain setting, where explanation only has an explanatory purpose. Despite their logical and coherent appearance favoring trust in the model Han et al. (2023), LLM generated CoT and self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model Agarwal et al. (2024), raising the need to accurately measure their faithfulness Jacovi and Goldberg (2020). This paper focuses on the case of measuring self-NLE faithfulness of LLM (i.e. in a predict-then-explain setting).
Numerous methods have been proposed to address this issue (see e.g. Lyu et al. (2024)). These approaches suffer from two main limitations: (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes Atanasova et al. (2023); Siegel et al. (2024, 2025); Matton et al. (2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without analyzing the actual neural activity underlying the prediction Wiegreffe et al. (2021); Parcalabescu and Frank (2024); Yeo et al. (2024).
@@ 2.1 Why Measure Explanation Faithfulness?
In their foundational paper Jacovi and Goldberg (2020) define model explanation faithfulness as "how accurately the explanation reflects the true reasoning process of the model", which is the definition generally adopted in the literature Lyu et al. (2024) and which we likewise follow throughout this work. Explanation faithfulness is commonly conflated with other explanation qualities such as readability and plausibility. This conflation can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair Luo et al. (2022) or harmful Kayser et al. (2024) decisions.
This tension between plausibility and faithfulness has become more pronounced with recent advances in NLP Agarwal et al. (2024). Unlike their non-generative predecessors, modern LLM are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations Rawte et al. (2023). The unstructured nature of these self-NLE pose greater challenges for faithfulness assessment compared to more constrained explainability techniques like attribution Wiegreffe et al. (2021) or counterfactual methods Madsen et al. (2024). As the use of LLM is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting underpins the urgent to evaluate self-NLE faithfulness.
@@ 2.2 How to Measure Self-NLE Faithfulness of LLM?
Despite the critical importance of self-NLE faithfulness, there is no consensus on how to quantitatively measure it Lyu et al. (2024). Numerous approaches have been proposed that we group in 3 categories.
@@
NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE Atanasova et al. (2023). Counterfactual Intervention (CI) methods employ auxiliary models to generate counterfactual texts designed to change the LLM outcome. The LLM is then prompted to produce a self-NLE to justify its new prediction. The self-NLE is deemed faithful if it aligns with the specific CI that caused the prediction change.
These CI-based approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE Atanasova et al. (2023); Siegel et al. (2024, 2025), and the granularity of the CI methods, whether at token or concept level Matton et al. (2025). However, these approaches face several limitations: (1) the CI may not be solely responsible for the change in prediction, as the model might base its new prediction based on another part of the input text after intervention, (2) CI methods overlook the model’s underlying reasoning process, which departs from the commonly adopted definition of explanation faithfulness.
@@
Another way to quantitatively measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE Parcalabescu and Frank (2024); Wiegreffe et al. (2021); Yeo et al. (2024). AA methods first compute attribution scores to interpret the model predictions and then measure how these scores correlate with other attribution scores computed to interpret related self-NLE. Higher correlation values indicate greater faithfulness in the model’s self-NLE.
AA approaches vary in their interpretation granularity (i.e., token or embedding level) and the post-hoc attribution method employed (gradient-based Sundararajan et al. (2017), SHAP Lundberg and Lee (2017) or activation patching Meng et al. (2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the neural activity within these blocks, leaving the model’s inner reasoning process only partially treated.
@@
The recent study by Tutek et al. (2025) emphasizes that CI methods actually measure contextual faithfulness by only perturbing the input text while keeping model parameters fixed. In contrast, it proposes to measure parametric faithfulness, which evaluates the faithfulness of self-NLE with respect to the model’s parameters. Their approach measures parametric faithfulness by assessing the effect of unlearning Liu et al. (2025) the parametric knowledge encoded in the reasoning steps of the self-NLE. The higher the change in prediction between the original model and the model having unlearned the reasoning steps, the more faithful the self-NLE. This approach inherently depends on the accuracy of the unlearning method employed, as either parameter localization or their erasure may be imprecise or inefficient for some target reasoning steps.
Most of the methods described above have been characterized as measuring self-consistency rather than true faithfulness Parcalabescu and Frank (2024), as they do not establish a direct link between the self-NLE and the model’s actual internal reasoning process. In the following, we propose a framework that improves upon these methods by explicitly examining the correspondence between self-NLE and interpretations of the model’s internal hidden states. However, we acknowledge our approach cannot perfectly capture the model’s complete inner reasoning process. Our approach therefore represents an advancement within the self-consistency paradigm by measuring consistency between neural internal activity and self-NLE. For convenience, in the following we keep on referring to faithfulness measurements.
@@ 3.2 Location
Transformer-based generative LLM can be viewed as a stack of decoder computational blocks Radford et al. (2018). The information flow through a single layer and from a layer to another can be described with the following equations:
@@
This characterization of the information flow through a Transformer block naturally highlights three possible locations, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information Elhage et al. (2021); Rai et al. (2024). These locations offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity Elhage et al. (2022). Thus, given a specific index/layer pair $(k,\ell)$ and following Equation 1, if $\mathcal{L}=\texttt{RS}$, $i^{\ell}_{k}(x)$ can be defined as the local neural interpretation of the hidden state $h^{\ell}_{k}$ on the residual stream.
@@ 3.3 Circuit
We restrict the interpretation of the hidden states to computational subgraphs that most significantly influences $f$ prediction by focusing on circuits Elhage et al. (2021). A circuit represents a sparse oriented subgraph of a neural network having an interpretable functional role, where nodes correspond to computational units (e.g. neuron, attention head, RS hidden state) and edges represent computation paths Wang et al. (2023); Räuker et al. (2023).
@@
A vast body of research has manually investigated task-specific circuits in Transformer-based LLM Bereska and Gavves (2024), such as entity resolution Wang et al. (2023), induction heads Olsson et al. (2022) or multi-hop reasoning Biran et al. (2024). Analyzing the faithfulness of an NLE for such tasks would consists in directly drawing inspiration from these established circuits to define $\mathcal{C}$.
@@
Alternatively, circuits can be obtained through automated methods, including techniques based on activation patching Meng et al. (2022); Syed et al. (2024); Conmy et al. (2023), backward attribution Ferrando et al. (2022); Ferrando and Voita (2024) or Transcoders Dunefsky et al. (2024); Ameisen et al. (2025). These automated approaches help address scenarios where circuits may not be well-documented in the literature.
@@ 3.4 Post-hoc Hidden State Interpreter
Recent research has shown that interpretable high-level features, referred to as concepts, appear to be linearly encoded within LLM latent space Elhage et al. (2022); Park et al. (2024). This computational understanding of concepts aligns with definitions from cognitive science Ruiz Luyten and van der Schaar (2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization Goguen (2005). This alignment makes concepts an ideal granularity level for model interpretation. Concept-based interpretations can be either obtained in a supervised or unsupervised manner, depending on the availability of texts with labeled concepts.
We make the general assumption that no such set of labeled concepts is available. Generating concept-based interpretations can be done by applying dictionary learning techniques Tošić and Frossard (2011) to project hidden states into a concept space. The semantics of the concepts are then determined once the concept dictionary has been learned. A popular approach for learning such dictionaries is to train sparse autoencoders (SAEHuben et al. (2023); Bricken et al. (2023) on the hidden states of $f$. Sparse concept activations are retrieved in the SAE hidden space to construct the concept-based activations as an interpretation.
@@
Another way of interpreting $f$ hidden states is to generate free text interpretations. Two recent and similar approaches, namely Selfie Chen et al. (2024) and Patchscopes Ghandeharioun et al. (2024), perform such interpretation generation by leveraging LLM decoding capabilities to produce natural language descriptions of hidden states. The interpretation is done by replacing the embedding of a placeholder token during a forward pass with the hidden state to be interpreted. The prompt of the interpretation forward pass is constructed to enhance free text interpretation expressivity.
@@ 3.5 Measuring Faithfulness
For $\mathcal{I}=\texttt{SAE}$, each interpretation in $i(x)$ is a set of vectors of concept activations with dimensions potentially exceeding a million Templeton et al. (2024), making direct comparison challenging. For a given interpretation $i^{\ell}_{k}(x)$, one way to simplify this comparison is to retrieve the concept labels with the highest activations. We formally define this simplification as applying the transformation $g(i^{\ell}_{k}(x))$ with $g:\mathbb{R}^{d}\rightarrow\mathcal{V}^{d^{\prime}}$ and $d^{\prime}\ll d$. The local faithfulness measure then evaluates whether the $d^{\prime}$ most salient concepts from $g(i^{\ell}_{k}(x))$ are present in $e(x)$.
For $\mathcal{I}=\texttt{Selfie}$, each interpretation in $i(x)$ consists of free-text interpretations. In this case, an interpretation $i^{\ell}_{k}(x)$ can be compared to $e(x)$ by assessing whether one logically corroborates the other. This comparison can be done either under an LLM-as-a-Judge settings Gu et al. (2024) or through human evaluation. We give a detailed example of how this comparison can be carried out in the case of 2-hop reasoning in Section 4.
@@ 4.1 Task Description
Multi-hop reasoning is a complex cognitive task in NLP that requires connecting a sequence of information to reach a conclusion Mavi et al. (2024). It consists in several single-hop operations Trivedi et al. (2022), which can individually be defined as triplets $(o_{1},r,o_{2})$ where $o_{1}$ is a source object, $r$ is a relation and $o_{2}$ is a target object. For example, the 2-hop reasoning statement "The country of origin of the movie maker that directed the movie Persona is Sweden" requires sequentially solving the two single-hop operations: $(o_{1},r_{1},o_{2})$ and $(o_{2},r_{2},o_{3})$ where $o_{1}$ is "Persona" (the movie), $r_{1}$ is "movie direction", $o_{2}$ is "Ingmar Bergman", $r_{2}$ is "country of origin" and the final answer $o_{3}$ is "Sweden". In this example, $o_{2}$ ("Ingmar Bergman") serves as the bridge object, connecting two reasoning operations and representing the intermediate step necessary to generate the correct final answer.
@@ 4.3 Measuring the Correctness of Self-NLE, Interpretation and Latent 2-hop Reasoning
This characterization of latent 2-hop reasoning is consistent with the one already mentioned in previous work Yang et al. (2024); Biran et al. (2024).
@@ 4.4 Highlighting the Tension Between Correctness and Faithfulness of 2-hop Self-NLE
  • ($C_{4}$) $f$ being deceptive. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $o_{2}\in i^{\ell}_{k}(x)$, $f$ appears to correctly perform latent reasoning but does not faithfully express this in its self-NLE, this way doing deceptive reasoning Greenblatt et al. (2024). This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while knowing the ground truth bridge object, raising a problem in $f$ alignment.
  • ($C_{5}$) $f$ doing shortcut learning. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\mathcal{C_{V}},o_{2}\notin i^{\ell}_{k}(x)$, the prediction is correct while $f$ seems to neither correctly perform latent reasoning nor generate a correct and faithful self-NLE, suggesting a potential case of shortcut learning Geirhos et al. (2020). This case is different from $C_{2}$ since shortcut learning implies correctly predicting $o_{3}$ directly based on $o_{1}$, whereas when $f$ is biased, the $o_{3}$ prediction is based on an incorrect bridge object $o_{2}$.

### v2

@@ 1 Introduction
Autoregressive Transformer-based (Vaswani et al., 2017) Large Language Models (LLM) have demonstrated striking performance in a wide range of Natural Language Processing (NLP) tasks (Chang et al., 2024). This achievement was largely attributable to the reasoning LLM abilities, enabling them to break down complex multi-step problems with Chain-of-Thought (CoT) (Wei et al., 2024) and generate self free text Natural Language Explanation (self-NLE) to support their answers (Wiegreffe and Marasovic, ; Huang et al., 2023). Generating CoT consists in prompting the LLM to generate an explanation in an explain-then-predict fashion, intending to improve the model’s performance. Conversely, generating self-NLE follows a predict-then-explain setting, where explanation only has an explanatory purpose. Despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM generated CoT and self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model (Agarwal et al., 2024), raising the need to accurately measure their faithfulness (Jacovi and Goldberg, 2020). This paper focuses on the case of measuring self-NLE faithfulness of LLM (i.e. in a predict-then-explain setting).
Numerous methods have been proposed to address this issue (see e.g. Lyu et al. (2024)). These approaches suffer from two main limitations: (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without analyzing the actual neural activity underlying the prediction (Wiegreffe et al., 2021; Parcalabescu and Frank, 2024; Yeo et al., 2024).
@@ 2.1 Why Measure Explanation Faithfulness?
In their foundational paper Jacovi and Goldberg (2020) define model explanation faithfulness as "how accurately the explanation reflects the true reasoning process of the model", which is the definition generally adopted in the literature (Lyu et al., 2024) and which we likewise follow throughout this work. Explanation faithfulness is commonly conflated with other explanation qualities such as readability and plausibility. This conflation can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions.
This tension between plausibility and faithfulness has become more pronounced with recent advances in NLP (Agarwal et al., 2024). Unlike their non-generative predecessors, modern LLM are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Rawte et al., 2023). The unstructured nature of these self-NLE pose greater challenges for faithfulness assessment compared to more constrained explainability techniques like attribution (Wiegreffe et al., 2021) or counterfactual methods (Madsen et al., 2024). As the use of LLM is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting underpins the urgent to evaluate self-NLE faithfulness.
@@ 2.2 How to Measure Self-NLE Faithfulness of LLM?
Despite the critical importance of self-NLE faithfulness, there is no consensus on how to quantitatively measure it (Lyu et al., 2024). Numerous approaches have been proposed that we group in 3 categories.
@@
NLE faithfulness can be assessed through behavioral tests that measure how perturbations in the input text affect both predictions and self-NLE (Atanasova et al., 2023). Counterfactual Intervention (CI) methods employ auxiliary models to generate counterfactual texts designed to change the LLM outcome. The LLM is then prompted to produce a self-NLE to justify its new prediction. The self-NLE is deemed faithful if it aligns with the specific CI that caused the prediction change.
These CI-based approaches mostly differ in two ways: how they measure consistency between the intervention and the resulting self-NLE (Atanasova et al., 2023; Siegel et al., 2024, 2025), and the granularity of the CI methods, whether at token or concept level (Matton et al., 2025). However, these approaches face several limitations: (1) the CI may not be solely responsible for the change in prediction, as the model might base its new prediction based on another part of the input text after intervention, (2) CI methods overlook the model’s underlying reasoning process, which departs from the commonly adopted definition of explanation faithfulness.
@@
Another way to quantitatively measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu and Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2024). AA methods first compute attribution scores to interpret the model predictions and then measure how these scores correlate with other attribution scores computed to interpret related self-NLE. Higher correlation values indicate greater faithfulness in the model’s self-NLE.
AA approaches vary in their interpretation granularity (i.e., token or embedding level) and the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg and Lee, 2017) or activation patching (Meng et al., 2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the neural activity within these blocks, leaving the model’s inner reasoning process only partially treated.
@@
The recent study by Tutek et al. (2025) emphasizes that CI methods actually measure contextual faithfulness by only perturbing the input text while keeping model parameters fixed. In contrast, it proposes to measure parametric faithfulness, which evaluates the faithfulness of self-NLE with respect to the model’s parameters. Their approach measures parametric faithfulness by assessing the effect of unlearning (Liu et al., 2025) the parametric knowledge encoded in the reasoning steps of the self-NLE. The higher the change in prediction between the original model and the model having unlearned the reasoning steps, the more faithful the self-NLE. This approach inherently depends on the accuracy of the unlearning method employed, as either parameter localization or their erasure may be imprecise or inefficient for some target reasoning steps.
Most of the methods described above have been characterized as measuring self-consistency rather than true faithfulness (Parcalabescu and Frank, 2024), as they do not establish a direct link between the self-NLE and the model’s actual internal reasoning process. In the following, we propose a framework that improves upon these methods by explicitly examining the correspondence between self-NLE and interpretations of the model’s internal hidden states. However, we acknowledge our approach cannot perfectly capture the model’s complete inner reasoning process. Our approach therefore represents an advancement within the self-consistency paradigm by measuring consistency between neural internal activity and self-NLE. For convenience, in the following we keep on referring to faithfulness measurements.
@@ 3.2 Location
Transformer-based generative LLM can be viewed as a stack of decoder computational blocks (Radford et al., 2018). The information flow through a single layer and from a layer to another can be described with the following equations:
@@
This characterization of the information flow through a Transformer block naturally highlights three possible locations, each offering different perspectives on the model’s information processing: the residual stream (RS) focusing on $h_{k}^{\ell}$, which serves as the main information pathway through the network; the multi-head attention (MHA) focusing on $a_{k}^{\ell}$, which captures token-to-token interactions; and the multi-layer perceptron (MLP) focusing on $m_{k}^{\ell}$ and performing non-linear transformations of the information (Elhage et al., 2021; Rai et al., 2024). These locations offer a relevant granularity to decode the internal activity of an LLM, since focusing on neurons alone can rarely be done due to their polysemanticity (Elhage et al., 2022). Thus, given a specific index/layer pair $(k,\ell)$ and following Equation 1, if $\mathcal{L}=\texttt{RS}$, $i^{\ell}_{k}(x)$ can be defined as the local neural interpretation of the hidden state $h^{\ell}_{k}$ on the residual stream.
@@ 3.3 Circuit
We restrict the interpretation of the hidden states to computational subgraphs that most significantly influences $f$ prediction by focusing on circuits (Elhage et al., 2021). A circuit represents a sparse oriented subgraph of a neural network having an interpretable functional role, where nodes correspond to computational units (e.g. neuron, attention head, RS hidden state) and edges represent computation paths (Wang et al., 2023; Räuker et al., 2023).
@@
A vast body of research has manually investigated task-specific circuits in Transformer-based LLM (Bereska and Gavves, 2024), such as entity resolution (Wang et al., 2023), induction heads (Olsson et al., 2022) or multi-hop reasoning (Biran et al., 2024). Analyzing the faithfulness of an NLE for such tasks would consists in directly drawing inspiration from these established circuits to define $\mathcal{C}$.
@@
Alternatively, circuits can be obtained through automated methods, including techniques based on activation patching (Meng et al., 2022; Syed et al., 2024; Conmy et al., 2023), backward attribution (Ferrando et al., 2022; Ferrando and Voita, 2024) or Transcoders (Dunefsky et al., 2024; Ameisen et al., 2025). These automated approaches help address scenarios where circuits may not be well-documented in the literature.
@@ 3.4 Post-hoc Hidden State Interpreter
Recent research has shown that interpretable high-level features, referred to as concepts, appear to be linearly encoded within LLM latent space (Elhage et al., 2022; Park et al., 2024). This computational understanding of concepts aligns with definitions from cognitive science (Ruiz Luyten and van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005). This alignment makes concepts an ideal granularity level for model interpretation. Concept-based interpretations can be either obtained in a supervised or unsupervised manner, depending on the availability of texts with labeled concepts.
We make the general assumption that no such set of labeled concepts is available. Generating concept-based interpretations can be done by applying dictionary learning techniques (Tošić and Frossard, 2011) to project hidden states into a concept space. The semantics of the concepts are then determined once the concept dictionary has been learned. A popular approach for learning such dictionaries is to train sparse autoencoders (SAE(Huben et al., 2023; Bricken et al., 2023) on the hidden states of $f$. Sparse concept activations are retrieved in the SAE hidden space to construct the concept-based activations as an interpretation.
@@
Another way of interpreting $f$ hidden states is to generate free text interpretations. Two recent and similar approaches, namely Selfie (Chen et al., 2024) and Patchscopes (Ghandeharioun et al., 2024), perform such interpretation generation by leveraging LLM decoding capabilities to produce natural language descriptions of hidden states. The interpretation is done by replacing the embedding of a placeholder token during a forward pass with the hidden state to be interpreted. The prompt of the interpretation forward pass is constructed to enhance free text interpretation expressivity.
@@ 3.5 Measuring Faithfulness
For $\mathcal{I}=\texttt{SAE}$, each interpretation in $i(x)$ is a set of vectors of concept activations with dimensions potentially exceeding a million (Templeton et al., 2024), making direct comparison challenging. For a given interpretation $i^{\ell}_{k}(x)$, one way to simplify this comparison is to retrieve the concept labels with the highest activations. We formally define this simplification as applying the transformation $g(i^{\ell}_{k}(x))$ with $g:\mathbb{R}^{d}\rightarrow\mathcal{V}^{d^{\prime}}$ and $d^{\prime}\ll d$. The local faithfulness measure then evaluates whether the $d^{\prime}$ most salient concepts from $g(i^{\ell}_{k}(x))$ are present in $e(x)$.
For $\mathcal{I}=\texttt{Selfie}$, each interpretation in $i(x)$ consists of free-text interpretations. In this case, an interpretation $i^{\ell}_{k}(x)$ can be compared to $e(x)$ by assessing whether one logically corroborates the other. This comparison can be done either under an LLM-as-a-Judge settings (Gu et al., 2024) or through human evaluation. We give a detailed example of how this comparison can be carried out in the case of 2-hop reasoning in Section 4.
@@ 4.1 Task Description
Multi-hop reasoning is a complex cognitive task in NLP that requires connecting a sequence of information to reach a conclusion (Mavi et al., 2024). It consists in several single-hop operations (Trivedi et al., 2022), which can individually be defined as triplets $(o_{1},r,o_{2})$ where $o_{1}$ is a source object, $r$ is a relation and $o_{2}$ is a target object. For example, the 2-hop reasoning statement "The country of origin of the movie maker that directed the movie Persona is Sweden" requires sequentially solving the two single-hop operations: $(o_{1},r_{1},o_{2})$ and $(o_{2},r_{2},o_{3})$ where $o_{1}$ is "Persona" (the movie), $r_{1}$ is "movie direction", $o_{2}$ is "Ingmar Bergman", $r_{2}$ is "country of origin" and the final answer $o_{3}$ is "Sweden". In this example, $o_{2}$ ("Ingmar Bergman") serves as the bridge object, connecting two reasoning operations and representing the intermediate step necessary to generate the correct final answer.
@@ 4.3 Measuring the Correctness of Self-NLE, Interpretation and Latent 2-hop Reasoning
This characterization of latent 2-hop reasoning is consistent with the one already mentioned in previous work (Yang et al., 2024; Biran et al., 2024).
@@ 4.4 Highlighting the Tension Between Correctness and Faithfulness of 2-hop Self-NLE
  • ($C_{4}$) $f$ being deceptive. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $o_{2}\in i^{\ell}_{k}(x)$, $f$ appears to correctly perform latent reasoning but does not faithfully express this in its self-NLE, this way doing deceptive reasoning (Greenblatt et al., 2024). This case is expected to be rare, otherwise highlighting a case where $f$ is not honest in its self-NLE while knowing the ground truth bridge object, raising a problem in $f$ alignment.
  • ($C_{5}$) $f$ doing shortcut learning. If $\widehat{o_{3}}=o_{3}$, $F(x,e,i)=0$, $\widehat{o_{2}}\neq o_{2}$ and $\forall(k,\ell)\in\mathcal{C_{V}},o_{2}\notin i^{\ell}_{k}(x)$, the prediction is correct while $f$ seems to neither correctly perform latent reasoning nor generate a correct and faithful self-NLE, suggesting a potential case of shortcut learning (Geirhos et al., 2020). This case is different from $C_{2}$ since shortcut learning implies correctly predicting $o_{3}$ directly based on $o_{1}$, whereas when $f$ is biased, the $o_{3}$ prediction is based on an incorrect bridge object $o_{2}$.